OpenStudy (anonymous):

write an equation in standard form with a slope of 2 and a y-intercept of 3

OpenStudy (anonymous):

use y=mx+b form where m is the slope and b is the y intercept

OpenStudy (anonymous):

y=2x+3

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

how do i make that into standard form?

OpenStudy (anonymous):

standard form is Ax + By = C get the x and y alone on one side

OpenStudy (anonymous):

2x+y=3?

OpenStudy (anonymous):

I think it's y-2x=3 give me a minute

OpenStudy (anonymous):

then we multiply both sides by -1 2x-y=-3

OpenStudy (anonymous):

Thank you